A total of $5,000 was invested, part of it at 5 % interest and the remainder at 7 %. If the total yearly interest amount is $325 , how much was invested at 5 %?

Answer by mananth(16946)

Part I 5.00% per annum ---x
part II 7.00% per annum ---y

x+y=5000------------------------1
5.00% x + 7.00% y= 325
Multiply by 100
5x+7y=32500.00 --------2
Multiply (1) by -5
we get
-5x-5y=-25000.00
Add this to (2)
2y=7500
divide by 2
y = 3750 Part I @ 7.00%
Balance 1250 part II @ 5.00%
CHECK
3750 --------- 7.00% ------- 262.50
1250 ------- 5.00% ------- 62.50
Total -------------- 325.00